# Fight like a Woman:

### v2

@@ Introduction
Since the results in Table 3 may capture differences between male and female judges beyond their gendered perspectives on domestic violence, we also estimate Equation (2) to compare the gender conviction rate gap for domestic violence cases against the gap for similar cases. Table 4 shows the estimated results associated with Equation (2). The first three columns use the sample of misdemeanors, therefore comparing crimes with similar sentence lengths to those in the domestic violence sub-sample. The last three columns use the sample of physical assault cases, which are similar in nature to domestic violence cases as they entail physical harm to another person.
@@
  • Notes: This table presents the estimated results of Equation (2). The first three columns compare differences in male and female judge conviction rates in domestic violence cases to the difference in conviction rates for misdemeanors. The last three columns compare differences in male and female judge conviction rates in domestic violence cases to the difference in conviction rates for other physical assault crimes. CR stands for Conviction Rate and DV stands for Domestic Violence.
Table 4 shows that the estimated coefficient on the female dummy ($\beta_{1}$) is small for the misdemeanors sample and close to zero for the physical assault sample. This suggests that the gender conviction rate gap is small or nonexistent for comparable crimes that are not classified as involving domestic violence.
More importantly, Table 4 shows that the coefficient on the interaction of judge gender and domestic violence ($\beta_{3}$) is statistically significant and positive in all regressions, regardless of included control variables. The effect is particularly strong for the physical assault sample (Columns (4)-(6)). This result suggests that the difference in conviction rates verified in Table 3 is neither fully explained by the severity or the nature of domestic violence offenses nor by female judges being tougher than male judges across the board.
The interaction coefficient captures the part of the gender conviction rate gap that is specific to domestic violence cases and is likely due to different approaches to domestic violence cases between male and female judges. For this reason, this result suggests the existence of a type of in-group bias according to the definition given by Shayo and Zussman (2011). [^10] In Appendix C, we propose a simple economic model formalizing the definition of in-group bias. In the next section, we explore the driving forces behind this in-group bias.
Understanding the Gender Conviction Rate Gap
@@
The first is suggestive evidence that the gap is explained by judges’ identification with their own gender (Section 5.1). We find that female judges are more likely than male judges to convict and simultaneously flag the case as an intimate partner violence offense, while female and male judges are equally likely to convict and not flag the case as an intimate partner violence offense. This result suggests that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4 because violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
@@
The second finding, discussed in Section 5.2, is that men and women respond differently to evidence presented in the same case. We present evidence that female judges are more likely than male judges to convict in cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating. This result suggests that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.
@@
In Appendix C, we propose a simple threshold-crossing model that formalizes the concepts of the representational and informational accounts.
Female Identity as a Driver for Conviction Rates
@@
In this section, we show that the gender of the judge matters more in cases where patriarchal social norms are particularly relevant: intimate partner violence. We interpret this finding as indicative that a major reason for the gender conviction gap is that domestic violence cases make gender identity more salient to judges, suggesting that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Section 4.
@@
We argue that domestic violence cases make gender identity more salient because sexist social norms and gender roles significantly affect the occurrence and punishment of intimate partner violence in multiple contexts. For instance, González and Rodríguez-Planas (2020) show that, among immigrant families, the incidence of intimate partner violence is positively affected by gender norms in the country of origin. Heise and Kotsadam (2015) show that the incidence of intimate partner violence in different countries is strongly associated with patriarchal social norms, such as male authority over women, justifications for wife beating, and laws and practices that disadvantage women in property ownership and other economic resources. Closer to our context, Perova and Reynolds (2017) show that police stations that specialize in domestic violence with specially trained officers significantly reduce the number of deaths of young women by intimate partner violence in metropolitan municipalities in Brazil. In terms of judicial decision-making on domestic violence, the negative stereotypes mentioned by Kafka et al. (2019) and Human Rights Watch (1991) concern female intimate partners.
@@
We, then, test whether female judges are more likely to make a conviction decision in a domestic violence case and flag the case as being related to intimate partner violence after controlling for case characteristics. If there is a gender conviction gap in cases flagged as intimate partner violence but not in cases without such a flag, then a stronger identification with the victim may be driving the gender lenience gap found in Section 4. If the gender conviction rate gap is present regardless of explicit mentions of intimate partner violence, then there must be other forces driving this lenience gap.
To implement this test, we interact the indicator for assault between intimate partners with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
Table 5 shows the results of these regressions. In Column (1), the outcome is a dummy variable indicating whether the case included intimate partner violence and resulted in a conviction. In Column (2), the outcome is a dummy variable indicating whether the case did not include intimate partner violence and also resulted in a conviction. In Column (3), the outcome is a dummy variable indicating whether the case resulted in a conviction, replicating the results from Table 3 for convenience.
We find that the probability of an intimate partner violence case resulting in conviction is 37.5% (9 p.p.) higher when the judge is female (Column (1)). However, the probability of a case not involving intimate partner violence resulting in conviction is the same regardless of the gender of the judge (Column (2)). These results suggest that the gender conviction rate gap is higher in cases where a female judge may identify more strongly with the victim compared to a male judge.
@@
  • The regressions in this table refer to outcomes for the sample of domestic violence crimes. The outcome in Column (1) is whether the case involved intimate partner violence and led to a conviction. The outcome in Column (2) is whether the case did not involve intimate partner violence and led to a conviction. The outcome in Column (3) is whether the case led to a conviction. All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
These findings are consistent with existing evidence that, in many contexts, judges make different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Many of these papers show that the differences intensify as group identity becomes more salient. This intensification may result from of a heightened opposition between two or more social groups (“us” versus “them”). For example, Shayo and Zussman (2011) show that Jewish Israeli judges in small claim courts favor Jewish litigants more intensely when they are more exposed to conflict-related terrorism. Closer to our case, Cai et al. (2022) show that judges analyzing divorce cases in China favor plaintiffs of their own gender more in areas where social norms are more conservative. In our case, we show that the gender gap is much more intense in cases where the form of domestic violence is more stereotypical. [^12] Note that most domestic violence cases against women are committed by partners and ex-partners (United Nations Office on Drugs and Crime, 2023). Consequently, the representational account (Boyd, Epstein, and Martin, 2010) is a plausible driving forcing of the in-group bias documented in Section 4.
@@
In this section, we show that female judges are more likely than male judges to convict in domestic violence cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating. We interpret this result as suggesting that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.
@@
Our dataset allows us to investigate whether the gender conviction rate gap is stronger in cases where the evidence set is less clearly incriminating. For each case, we observe whether the defendant was arrested because they were caught in the act of committing the crime (i.e., in flagrante delicto). These cases have more objective evidence, often including testimony from police officers who witnessed the offense.
If the gender conviction rate gap is stronger in cases where the defendant was not caught red-handed, then this may be because female and male judges interpret more subjective pieces of evidence differently (informational account). If the gender conviction rate gap is similar across domestic violence cases regardless of the type of evidence presented, then there must be other forces driving this lenience gap.
To measure the gender conviction rate gap for cases with different evidence types, we re-estimate Equations (1) and (2) using two sub-samples. The first restricts our sample to cases where the defendant was caught red-handed, while the second is restricted to cases where the defendant was not caught in flagrante delicto.
Table 6 reports regression results for the two different samples. The odd columns include cases where the defendant was caught red-handed, while the even columns include cases where this objective evidence was not presented. Columns (1) and (2) report the results associated with Equation (1), while Columns (3)-(6) report the results associated with the interaction model described in Equation (2).
@@
  • Notes: This table compares the determinants of the probability of conviction. Cases are split into a sub-sample in which the defendant was caught red-handed and a sample in which the defendant was not, requiring the justice system to rely on less objective evidence. The first two columns are constructed using the sample of domestic violence cases. Columns 3 and 4 use the sample of misdemeanors. Columns 5 and 6 use the sample of physical assault crimes. CR stands for Conviction Rate and DV stands for Domestic Violence.
@@
We find that a gender conviction rate gap exists only in those cases where the defendant was not caught red-handed. We only find significant results for the female coefficient ($\beta$) in Column (2) and the interaction coefficient ($\beta_{3}$) in Column (6). These results suggest that female and male judges interpret less objective evidence differently, while they interpret objective evidence (cases flagged as in flagrante delicto) in similar ways. These different interpretations of more subjective cases are consistent with the informational account being one possible driving force of the in-group bias documented in Section 4.
Discussion: Career Incentives and a Potential Underestimation of the Gap
In this section, we analyze the association between the strength of career incentives and the size of the conviction gender gap. We show that male and female judges differ more intensely at the end of their careers. This corresponds with the period where they face the lowest probability of upcoming performance evaluation for promotion to the Appeals Court. [^13] See Appendix B for more information on the proportion of cases analyzed by judges who apply for promotion in the near future Career pressures may lead early-career male and female judges to act more similarly, possibly in response to feedback from Appeals Court magistrates, who are overwhelmingly male (Laneuville, 2024).
Since Brazilian judges cannot be fired and must receive the same salary, promotion is the main career incentive faced by magistrates in our sample. Judges are unelected officials who may occupy four positions during their careers in the Trial Court (substitute, initial, intermediary, and final). At the final stage of their careers, they may be promoted to the Appeals Court. To be promoted to the next career stage, Trial Court Judges must apply for promotion opportunities and be approved by a vote of the Appeals Court Judges. Since very few judges are promoted to the Appeals Court (Laneuville, 2024), career incentives are stronger for trial judges who are not at the final stage of the Trial Court. [^14] Appendix B provides descriptive statistics showing that final-stage judges face weaker career incentives.
Since adopting different sentencing criteria from peer judges may hamper chances for promotion, we investigate whether judges who face weak career incentives are more likely to differ from their colleagues, and whether these differences correspond with judge gender. If we observe a larger gender conviction gap for final-stage judges, then weaker career incentives may explain the gender leniency gap in domestic violence cases documented in Section 4. If judges in all career stages have similar gender conviction rates, then there must be other forces driving this lenience gap.
To measure the gender conviction gap for judges in different career stages, we re-estimate Equations (1) and (2) using two sub-samples. The first restricts our sample to cases assigned to final-stage judges, while the second restricts the sample to cases assigned to judges in earlier stages of the Trial Court.
Table 7 presents the estimated coefficients of Equations (1) and (2) for these two sub-samples. Results for the sample of final-stage judges are provided in the odd columns. Results for the sample of early-stage judges are provided in the even columns. The header indicates whether the sample contains only domestic violence cases, misdemeanor cases, or physical assault cases.
Table 7: Effect of Judge Gender on the Probability of Conviction by Judge Career Stage
Sample: Sample: Sample:
Domestic Violence Misdemeanors Physical Assault
Final Early Final Early Final Early
@@
  • Notes: This table shows results of regressing the conviction indicator on the indicator variables for female judges, domestic violence cases, and the interaction female x domestic violence for six different samples. Column (1) shows the results for the domestic violence cases assigned to final-stage judges. Column (2) shows the results for domestic violence cases assigned to early-stage judges. Column (3) shows the results for misdemeanor cases assigned to final-stage judges. Column (4) shows the results for misdemeanor cases assigned to early-stage judges. Column (5) shows the results for physical assault cases assigned to final-stage judges. Column (6) shows the results for cases assigned to early-stage judges. All regressions include quarter x court district fixed effects. CR stands for Conviction Rate and DV stands for Domestic Violence.
We find that the gender conviction rate gap is much stronger for final-stage judges than for early-stage judges. In the sample of domestic violence cases, final-stage female judges are almost twice as likely to convict as their male peers (23 p.p. larger), while early-stage judges have similar conviction rates across genders. We reach qualitatively similar conclusions when analyzing the interaction coefficients in the misdemeanors and other physical assault samples. Hence, the bulk of the gender gap discussed in this paper comes from judges at the end of their careers who face lower pressure to perform. If career pressures affect judges’ verdicts, then early-career judges of different genders may have different views on domestic violence than their conviction rates might suggest.

### v3

@@ Introduction
Since the results in Table 3 may capture differences between male and female judges beyond their gendered perspectives on domestic violence, we also estimate Equation (2) to compare the gender conviction-rate gap for domestic violence cases against this gap for similar cases. Table 4 shows the estimated results associated with Equation (2). The first three columns use the sample of misdemeanors, therefore comparing crimes with similar sentence lengths to those in the domestic violence sub-sample. The last three columns use the sample of physical assault cases, which are similar in nature to domestic violence cases as they entail physical harm to another person.
Table 4 shows that the estimated coefficient on the female dummy ($\beta_{1}$) is small for the misdemeanors sample and close to zero for the physical assault sample. This suggests that the gender conviction-rate gap is small or nonexistent for comparable crimes that are not classified as involving domestic violence.
More importantly, Table 4 shows that the coefficient on the interaction of judge gender and domestic violence ($\beta_{3}$) is statistically significant and positive in all regressions, regardless of included control variables. The effect is particularly strong for the physical assault sample (Columns (4)-(6)). This result suggests that the gender difference in conviction rates verified in Table 3 is neither fully explained by the severity or the nature of domestic violence offenses nor by female judges being stricter than male judges in all rulings.
Consequently, the interaction coefficient captures the part of the gender conviction-rate gap that is specific to domestic violence cases and is likely due to different approaches to domestic violence cases between male and female judges. For this reason, this result suggests the existence of a type of in-group bias according to the definition given by Shayo and Zussman (2011). [^12] In Appendix B, we propose a simple economic model formalizing the definition of in-group bias. In the next section, we explore the driving forces behind this in-group bias, while, in Section 6, we investigate its consequences for the cases’ later stages in the Justice System and the defendants’ future behavior.
@@
  • Notes: This table presents the estimated results of Equation (2). The first three columns compare differences in male and female judge conviction rates in domestic violence cases to the gender difference in conviction rates for misdemeanors. The last three columns compare differences in male and female judge conviction rates in domestic violence cases to the gender difference in conviction rates for other physical assault crimes. CR stands for Conviction Rate and DV stands for Domestic Violence.
Understanding the Gender Conviction-Rate Gap
@@
The first finding is suggestive evidence that the gap is explained by judges’ identification with their own gender (Section 5.1). We find that female judges are more likely than male judges to convict and simultaneously flag the case as an intimate partner violence offense, while female and male judges are equally likely to convict and not flag the case as an intimate partner violence offense. This result suggests that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4 because violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient (Heise and Kotsadam, 2015; Tur-Prats, 2019; González and Rodríguez-Planas, 2020).
@@
The second finding, discussed in Section 5.2, is that men and women respond differently to evidence presented in the same case. We show that female judges are more likely than male judges to convict in cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating (i.e., when the defendant was caught red-handed). This result suggests that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.
@@
In Appendix B, we propose a simple threshold-crossing model that formalizes the concepts of the representational and informational accounts.
Female Identity as a Driver for Larger Conviction Rates
@@
In this section, we show that the gender of the judge matters more in cases where patriarchal social norms are particularly relevant: intimate partner violence. We interpret this finding as indicative that a major reason for the gender conviction-rate gap is that domestic violence cases make gender identity more salient to judges, suggesting that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Section 4 because female judges tend to protect member of their own group more intensely than male judges.
@@
We argue that domestic violence cases make gender identity more salient because sexist social norms and gender roles significantly affect the occurrence and punishment of intimate partner violence in multiple contexts. For instance, González and Rodríguez-Planas (2020) show that, among immigrant families, the incidence of intimate partner violence is positively affected by gender norms in the country of origin. Heise and Kotsadam (2015) show that the incidence of intimate partner violence in different countries is strongly associated with patriarchal social norms, such as male authority over women, justifications for wife beating, and laws and practices that disadvantage women in property ownership. Closer to our context, Perova and Reynolds (2017) show that police stations that specialize in domestic violence with specially trained officers significantly reduce the number of deaths of young women by intimate partner violence in metropolitan municipalities in Brazil. In terms of judicial decision-making on domestic violence, the negative stereotypes mentioned by Kafka et al. (2019) and Human Rights Watch (1991) concern female intimate partners.
@@
We, then, test whether female judges are more likely to make a conviction decision in a domestic violence case and simultaneously flag the case as being related to intimate partner violence after controlling for case characteristics. If there is a gender conviction-rate gap in cases flagged as intimate partner violence but not in cases without such a flag, then a stronger identification with the victim may be driving the gender lenience gap found in Section 4. If the gender conviction-rate gap is present regardless of explicit mentions to intimate partner violence, then there must be other forces driving this lenience gap.
To implement this test, we interact the “intimate partner violence” indicator with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
Table 5 shows the results of these regressions. In Column (1), the outcome is a dummy variable indicating whether the judge’s ruling mentioned intimate partner violence and reached a conviction decision. In Column (2), the outcome is a dummy variable indicating whether the judge’s ruling did not mention intimate partner violence but reached a conviction decision. In Column (3), the outcome is a dummy variable indicating whether the case resulted in a conviction, replicating the results from Table 3 for convenience.
We find that female judges are 37.5% (9 p.p.) more likely than male judges to convict the defendant and simultaneously flag the case as intimate partner violence (Column (1)). However, female judges are as likely as male judges to convict the defendant while not flagging the case as intimate partner violence (Column (2)). These results suggest that the gender conviction-rate gap is higher in cases where a female judge may identify more strongly with the victim compared to a male judge.
@@
  • The regressions in this table refer to outcomes for the sample of domestic violence crimes only. The outcome in Column (1) is whether the case involved intimate partner violence and led to a conviction. The outcome in Column (2) is whether the case did not involve intimate partner violence and led to a conviction. The outcome in Column (3) is whether the case led to a conviction. All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
These findings are consistent with existing evidence that, in many contexts, judges make different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Many of these papers show that the differences intensify as group identity becomes more salient. This intensification may result from of a heightened opposition between two or more social groups (“us” versus “them”). For example, Shayo and Zussman (2011) show that Jewish Israeli judges in small claim courts favor Jewish litigants more intensely when they are more exposed to conflict-related terrorism. Closer to our case, Cai et al. (2022) show that judges analyzing divorce cases in China favor plaintiffs of their own gender more in areas where social norms are more conservative. In our case, we show that the gender conviction-rate gap is much more intense in cases where the form of domestic violence is more stereotypical. [^14] Note that most domestic violence cases against women are committed by partners and ex-partners (United Nations Office on Drugs and Crime, 2023). Consequently, the representational account (Boyd, Epstein, and Martin, 2010) is a plausible driving forcing of the in-group bias documented in Section 4 because female judges tend to protect members of their own group more intensely than male judges.
@@
In this section, we show that female judges are more likely than male judges to convict in domestic violence cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating. We interpret this result as suggesting that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4 because female judges process the information contained in domestic violence cases differently from their male peers.
@@
Our dataset allows us to investigate whether the gender conviction-rate gap is stronger in cases where the evidence set is less clearly incriminating. For each case, we observe whether the defendant was arrested because they were caught in the act of committing the crime (i.e., in flagrante delicto). These cases have more objective evidence, often including testimony from police officers who directly witnessed the offense.
If the gender conviction-rate gap is stronger in cases where the defendant was not caught red-handed, then the gender leniency gap may be because female and male judges interpret more subjective pieces of evidence differently (informational account). If the gender conviction-rate gap is similar across domestic violence cases regardless of the type of evidence presented, then there must be other forces driving this lenience gap.
To measure the gender conviction-rate gap for cases with different evidence types, we re-estimate Equations (1) and (2) using two sub-samples. The first regression restricts our sample to cases where the defendant was caught red-handed, while the second regression restricts our sample to cases where the defendant was not caught in flagrante delicto.
Table 6 reports regression results for the two different samples. The odd columns include cases where the defendant was caught red-handed, while the even columns include cases where this objective evidence was not present. Columns (1) and (2) report the results associated with Equation (1), while Columns (3)-(6) report the results associated with the interaction model described in Equation (2).
@@
  • Notes: This table compares the determinants of the probability of conviction. Cases are split into a sub-sample in which the defendant was caught red-handed (odd columns) and a sample in which the defendant was not caught red-handed (even columns), requiring the justice system to rely on less objective evidence. The first two columns are constructed using the sample of domestic violence cases only. Columns (3) and (4) use the sample of misdemeanors. Columns (5) and (6) use the sample of physical assault crimes. CR stands for Conviction Rate and DV stands for Domestic Violence. All regressions include quarter-district fixed effects and a full set of controls for judge career stages.
@@
We find that a gender conviction-rate gap exists only in those cases where the defendant was not caught red-handed. We only find significant results for the female coefficient ($\beta$) in Column (2) and the interaction coefficient ($\beta_{3}$) in Column (6). These results suggest that female and male judges interpret less objective evidence differently, while they interpret objective evidence (cases flagged as in flagrante delicto) in similar ways. These different interpretations of more subjective cases are consistent with the informational account being one possible driving force of the in-group bias documented in Section 4.
Consequences of the Gender Conviction-Rate Gap
A gender conviction-rate gap for domestic violence cases, as found in Section 4, may have consequences for the cases’ later stages in the Justice System and the defendants’ future behavior. In particular, defense attorneys, district attorneys, Appeals Court judges, and defendants may change their behavior when facing a female judge in a domestic violence case.
In Section 6.1, we investigate whether stricter female judges may lead to more appeals and ruling reversals in domestic violence cases. Although we may expect positive gender appeal-rate and reversal-rate gaps due to the fact that stricter judges are more frequently reversed by the Appeals Court (Possebom, 2023, Figure 1), we find that female judges’ rulings in domestic violence cases have similar outcomes in the Appeals Court when compared to male judges’ rulings.
In Section 6.2, we investigate whether stricter female judges may lead to more recidivism by defendants in domestic violence cases. Since stricter female judges might be more likely to convict the type of defendants whose recidivism increases when punished (Acerenza, Possebom, and Sant’Anna, 2024), we may expect a positive gender recidivism gap. Despite this expectation, we find that female and male judges face similar recidivism probabilities in their domestic violence cases.
Consequences of the Gender Conviction-Rate Gap at the Appeals Court
A gender conviction-rate gap for domestic violence cases may impact how the defense attorneys, district attorneys and Appeals Court judges treat domestic violence cases after the Trial Judge has issued a ruling. In this section, we investigate whether these agents change their behavior when facing a female judge in a domestic violence case.
On the one hand, the phenomenon of stricter female judges for domestic violence cases may lead to more appeals and reversals of trial judges’ rulings. This positive gender gap may be expected for two reasons.
First, Possebom (2023, Figure 1 and Table J.1) finds that rulings by strict judges are more frequently reversed by the Appeals Court than rulings by more lenient judges, partially because defense attorneys are more likely to appeal than district attorneys. Combining this result with our finding that female judges are stricter than male judges when analyzing domestic violence cases, we may expect a positive gender gap in the probability of domestic violence cases being analyzed by the Appeals Court and in the probability of domestic violence rulings being reversed by the Appeals Court.
Second, Laneuville (2024, Table 3) states that 93.55% of Appeals Court judges who were active between 2007 and 2022 were men. If we assume that judges of both genders behave similarly at the Trial Court and the Appeals Court, our results suggest that male Appeals Court judges may be more lenient when analyzing domestic violence cases and reverse Trial Judges’ conviction rulings more frequently. Consequently, we may expect a positive gender gap in the probability of domestic violence rulings being reversed by the Appeals Court.
On the other hand, the informational account for the gender conviction-rate gap in domestic violence cases (Section 5.2) may lead to a smaller probability of female judges having their rulings reversed by the Appeals Court. According to the informational account of the in-group bias, female judges analyze domestic violence evidence in a different way, possibly implying that they write more complete or careful rulings. When we combine this possibility with the fact that the Appeals Court focuses more on the ruling’s form than the case’s substance, we may expect a negative gender gap in the probability of domestic violence rulings being reversed by the Appeals Court.
We investigate which of these two forces dominates by analyzing if there is a gender gap in the probability of domestic violence cases being analyzed by the Appeals Court and in the probability of domestic violence rulings being reversed by the Appeals Court. To do so, we re-estimate Equations (1) and (2) using two new outcome variables. First, our left-hand side variable is an indicator that equals 1 if there is an appeal process associated with the case and 0 otherwise. Second, our outcome variable is an indicator that equals 1 if the Appeals Court reversed the Trial Judge’s ruling and 0 otherwise. To avoid sample selection issues, note that the “zero” group includes all cases that were not analyzed by the Appeals Court.
Table 7 shows the estimated coefficients of Equations (1) and (2) when we use our appeal and reversal indicators as outcome variables. The left-hand side variable is the appeal indicator in the odd columns, while it is the reversal indicator in the even columns. The first two columns are constructed using the sample of domestic violence cases only (Equation (1)). Columns (3) and (4) use the sample of misdemeanors (Equation (2)), while Columns (5) and (6) use the sample of physical assault crimes (Equation (2)). All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
There are three results worth highlighting. First, the sign and significance of the female coefficient in Columns (3) and (5) show that female judges are slightly more likely to be appealed in misdemeanor and physical assault cases, respectively. This finding is not surprising because female judges are stricter than male judges (Column (3) in Table 4), and defense attorneys appeal more frequently than district attorneys (Possebom, 2023, Table J.1). Second, because female coefficients are statistically null in Columns (2), (4) and (6), we find that, even though female trial judges are more frequently appealed, there is no gender difference in the reversal probability. Third and more importantly, the coefficients associated with the consequences of the in-group bias (“female” in Columns (1) and (2) and “DV x Female” in Columns (3)-(6)) are not statistically significant. [^15] Remember that the in-group bias captures the part of the gender conviction-rate gap that is specific to domestic violence cases. The null result in the “female” coefficient in Columns (1) and (2) means that female and male trial judges are equally likely to be appealed and reversed in domestic violence cases. The “DV x Female” coefficient in Columns (3)-(6) means that the gender gap in the probability of being appealed and reversed is not significantly more intense in domestic violence cases than in other comparable cases.
These null results suggest that the forces driving the gender appeal-rate and reversal-rate gaps upwards and downwards compensate each other. In other words, even though female judges are stricter than male judges when analyzing domestic violence cases, their rulings are complete and careful enough so that they do not generate more reversal decisions by the Appeals Court.
Table 7: Effect of Judge Gender on the Probability of a Case Being Analyzed by the Appeals Court and of a Trial Sentence Being Reversed by the Appeals Court
Sample: DV only Sample: Misdemeanors Sample: Physical Assault
Appealed Reversed Appealed Reversed Appealed Reversed
@@
Career Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
  • Notes: This table shows the results of Equations (1) and (2) using two new outcome variables. In the odd columns, the outcome variable is binary and indicates whether the case was analyzed by the Appeals Court, while, in the even columns, the outcome variable is binary and indicates whether the trial judge’s sentence was reversed by the Appeals Court. The first two columns are constructed using the sample of domestic violence cases (Equation (1)). Columns (3) and (4) use the sample of misdemeanors (Equation (2)). Columns (5) and (6) use the sample of physical assault crimes (Equation (2)). DV stands for Domestic Violence. All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
Consequences of the Gender Conviction-Rate Gap for the Defendants’ Recidivism Behavior
A gender conviction-rate gap for domestic violence cases may impact how defendants behave after their sentences. In this section, we investigate whether these agents change their behavior when facing a female judge in a domestic violence case.
On the one hand, the phenomenon of stricter female judges for domestic violence cases may increase the probability of recidivism. This positive gender gap may be expected for two reasons.
First, Acerenza, Possebom, and Sant’Anna (2024) analyzes misdemeanor cases in São Paulo during our sample period and finds that punishment increases two-year recidivism for agents who would be punished only by strict judges (type A defendant) while it decreases two-year recidivism for agents who would be punished by most judges, including the lenient judges (type B defendant). In our context, we find that female judges are stricter than male judges (Table 4, Column (3)), especially for domestic violence cases (Table 4, Columns (3) and (6)). Moreover, if we assume that female judges are stricter because they have a harsher bar for conviction (instead of having more information about domestic violence cases), then we expect female judges to convict more agents who are more likely to recidivate when punished (type A defendant) while being similarly likely to convict type B defendants. Consequently, we may expect a positive gender gap in the probability of a defendant recidivating.
Second, domestic violence victims may be encouraged by conviction decisions to accuse their aggressors if they recidivate by committing a new domestic violence offense. If this is the case, we may expect a positive gender gap in the recidivism probability because female judges are stricter than male judges when analyzing domestic violence cases (Section 4). To avoid this mechanical increase in the recidivism probability, we also analyze the effect of a female judge on the probability of committing a new crime that is not a domestic violence offense.
On the other hand, the representational account for the gender conviction-rate gap in domestic violence cases (Section 5.1) may lead to a smaller probability of recidivism. For example, female judges may be able to directly influence the behavior of the defendants by better signaling that violence against women is not tolerated. If the typical defendant in a domestic violence case is more likely to recidivate in a similar offense than to commit other types of crime, female judges may lead to lower recidivism. Consequently, we may expect a negative gender gap in the recidivism probability.
We investigate which of these two forces dominates by analyzing if there is a gender gap in the probability of a defendant recidivating within two years of the case’s final sentence. To do so, we re-estimate Equations (1) and (2) using two new outcome variables. First, our left-hand side variable is an indicator that equals 1 if the defendant commits any new criminal offense within two years of their final sentence and 0 otherwise. Second, our outcome variable is an indicator that equals 1 if the defendant commits any new non-domestic violence offense within two years of their final sentence and 0 otherwise.
Table 8 shows the estimated coefficients of Equations (1) and (2) when we use these two measures of recidivism as outcome variables. The left-hand side variable is the any-recidivism indicator in the odd columns, while it is the non-domestic violence recidivism indicator in the even columns. The first two columns are constructed using the sample of domestic violence cases only (Equation (1)). Columns (3) and (4) use the sample of misdemeanors (Equation (2)), while Columns (5) and (6) use the sample of physical assault crimes (Equation (2)). All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
Note that the coefficients associated with the consequences of the in-group bias (“female” in Columns (1) and (2) and “DV x Female” in Columns (3)-(6)) are not statistically significant. [^16] Remember that the in-group bias captures the part of the gender conviction-rate gap that is specific to domestic violence cases. We also highlight that the relevant point estimates in Columns (1)-(4) are either small or negative.
These results suggest that the forces driving the gender recidivism gap upwards and downwards compensate each other. In other words, even though female judges are stricter than male judges when analyzing domestic violence cases, their stricter rulings are not criminogenic.
Table 8: Effect of Judge Gender on the Probability of a Defendant Recidivating
Sample: DV only Sample: Misdemeanors Sample: Physical Assault
Any Excluding DV Any Excluding DV Any Excluding DV
(1) (2) (3) (4) (5) (6)
Female $0.01$ $-0.02$ $0.01$ $0.00$ $-0.02$ $-0.04$
$(0.04)$ $(0.04)$ $(0.01)$ $(0.01)$ $(0.05)$ $(0.05)$
DV $-0.02$ $-0.06^{***}$ $0.02$ $-0.01$
$(0.02)$ $(0.01)$ $(0.04)$ $(0.04)$
DV x Female $0.01$ $0.00$ $0.04$ $0.03$
$(0.02)$ $(0.02)$ $(0.06)$ $(0.05)$
Men’s Average (Non-DV) $0.35$ $0.35$ $0.25$ $0.24$
Men’s Average (All) $0.32$ $0.28$ $0.35$ $0.34$ $0.30$ $0.27$
Career Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Num. obs. $2,665$ $2,665$ $21,319$ $21,319$ $3,467$ $3,467$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table shows the results of Equations (1) and (2) using two new outcome variables. In the odd columns, the outcome variable is binary and indicates whether the case’s defendant was prosecuted for any crime after their final sentence’s date, while, in the even columns, the outcome variable is binary and indicates whether the case’s defendant was prosecuted for any crime that is not a domestic violence offense. The first two columns are constructed using the sample of domestic violence cases (Equation (1)). Columns (3) and (4) use the sample of misdemeanors (Equation (2)). Columns (5) and (6) use the sample of physical assault crimes (Equation (2)). DV stands for Domestic Violence. All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.